Luke mixed two fruit juices to create a fruit punch.
• Luke bought 5 gallons of orange juice for a total cost of $15.
• The cost per gallon of pineapple juice is 35% more than the cost per gallon of orange juice.
• Luke bought enough pineapple juice that, when he mixed it with the orange juice, the fruit punch had a value of $3.16 per gallon.

Approximately what percent of the fruit punch was pineapple juice?

Answer by jorel555(1290)

5 gallons o.j.@ $15=$3/gallon
1.35*3=$4.06/gallon of pineapple juice
Let n be the amount of pineapple juice:
3(5)+4.06(n)=3.16(5+n)
15+4.06n=15.8+3.16n
.9n=.8
n=.8/.9=.8888 gallons of pineapple juice added
.88888/5.888888=15% pineapple juice in the fruit punch!!!!!!!